# Phase space analysis of CCDM cosmologies

### v1

@@ III.3 Model: $\Gamma=\Gamma_{0}H^{2}$
In this case, we have constructed the model considering the particle creation rate $\Gamma=\Gamma_{0}H^{2}$ where $\Gamma_{0}$ is a constant having dimension of the inverse of the Hubble rate. We shall also discuss how $\Gamma$ influences the universe’s evolution for both the fluids.
III.3.1 One fluid system
Following the similar arguments as in the earlier cases with one-fluid system, for this model, the one dimensional autonomous system can be reads as:
$z^{\prime}=\frac{3}{2}(1-z)\left(z-\frac{\beta(1-z)}{3}\right),$ (32)
where $\beta=\Gamma_{0}H_{0}$ is the dimensionless constant and it takes positive values. If we solve the quadratic equation $z^{\prime}=0$, we get two critical points, namely, $C^{\prime}_{0}(1,1)$ and $C^{\prime}_{1}\left(1,\frac{\beta}{3+\beta}\right)$, where first and second component of the critical points correspond to dark-matter density parameter and the value of $z$ variable respectively. The point $C^{\prime}_{0}$ is always stable and represents decelerating phase of the universe. Again, $C^{\prime}_{1}$ corresponds to an accelerating phase of the universe whenever $C^{\prime}_{1}$ exists, i.e. $\beta>0$, and this point is always unstable. The directions of the vector field defined in (32) are given in the upper left plot of the Fig. 5. In this case $q=\frac{1}{2}\left(1-\frac{\beta(1-z)}{z}\right)$. Then, the critical point $C^{\prime}_{1}$ corresponds to $q=-1$ that is, a de Sitter solution. At the present time, which corresponds to $z=1/2$, we have $q=\frac{1}{2}(1-\beta)$. Then, to have an acceleration at the present time we need $\beta>1$. On the other hand, for $\beta<3$ we have $\frac{\beta}{\beta+3}<\frac{1}{2}$. Thus, in this situation, the universe passes from an accelerating phase to a decelerating one at the late time. On the contrary, when $\beta>3$, we have $\frac{\beta}{\beta+3}>\frac{1}{2}$, and thus, the universe is always accelerating. Note that when $\beta=3$, the critical point coincides with $1/2$, this means at the present time the universe is in a de Sitter phase. In fact, when $31<\beta<$ 3, because in this case the universe starts in a de Sitter phase and at the present time it accelerates (finally, at late times it decelerates) always in a non-phantom phase. On the contrary, for $\beta>3$ the universe is always in a phantom phase. This one-fluid system can predict a transition of our universe from the present (or past) accelerating era to future decelerating era.
III.3.2 Two fluid system
@@
where $\beta=\Gamma_{0}H_{0}$ is a dimensionless and positive constant. The above system has singularity at $z=0$. We multiply to the right hand sides of (33) by $z$ to make the system regularised. After doing this, the regularized system becomes,
$x^{\prime}=$ $x(1-x)\left(3wz+\beta(1-z)\right),$ (34a)
$z^{\prime}=$ $\frac{3}{2}z(1-z)\left[\left(1+w(1-x)\right)z-\frac{\beta x(1-z)}{3}\right].$ (34b)
From eqn. (24), it is clear that our physical domain is $R_{3}=[0,1]^{2}$ and the dynamical system (34) gives $x=0,~{}x=1,~{}z=0,~{}z=1$ as the invariant manifolds which lead the domain $R_{3}$ to being a positively invariant set. Being $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\beta x(1-z)}{z}\right]$ in this case, the requirement for realizing an accelerating phase at the present moment (i.e. $z=1/2$), the physical trajectories should satisfy the relation $(1-\beta)x+(1+3w)(1-x)<0$. In the following we describe the qualitative features of the above dynamical system for different values of $w$.
One fluid system
Critical point Existence Stability $\Omega_{\rm dm}$ $q$ Acceleration
$C^{\prime}_{0}(1,1)$ Always Always Stable 1 $\frac{1}{2}$ No
$C^{\prime}_{1}\left(1,\frac{\beta}{3+\beta}\right)$ $\beta>0$ Always Unstable 1 $-1$ Yes
Two fluid system
Critical point Existence Eigenvalue Stability $\Omega_{f}$ $\Omega_{\rm dm}$ $q$ Acceleration
$C_{0}(0,0)$ Always $(\beta,0)$ Unstable if $w>-1$; 1 0 Undefined Undetermined
Non-hyperbolic Saddle if $w<-1$
$C_{1}(1,0)$ Always $\left(-\beta,-\frac{\beta}{2}\right)$ Stable 0 1 $\infty$- ∞ Yes
$C_{2}(0,1)$ Always $\left(3w,-\frac{3}{2}(1+w)\right)$ Stable if $-1<w<0$; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if $w<-1$ and if $w>0$
@@
$C_{4}\left(1,\frac{\beta}{\beta+3}\right)$ $\beta>0$ $\left(-\frac{3\beta(1+w)}{3+\beta},\frac{3\beta}{2(3+\beta)}\right)$ Saddle if $w>-1$ 0 1 $-1$ Yes
Unstable if $w<-1$
$C_{5}\left(0,z_{c}\right)$ $w=-1$ $\left(\beta-(\beta+3)z_{c},0\right)$ Stable if $z_{c}>\frac{\beta}{\beta+3}$ $1$ 00 $-1$ Yes
@@
$C_{6}\left(x_{c},\frac{\beta}{\beta+3}\right)$ $w=-1$ $\left(0,\frac{3\beta x_{c}}{2(3+\beta)}\right)$ Unstable $1-x_{c}$ $x_{c}$ $-1$ Yes
@@
Table 3: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (32) and (34) with the matter creation rate $\Gamma=\Gamma_{0}H^{2}$.
@@
Figure 5: Description of the phase space of the one-fluid and two-fluid systems for the matter creation rate $\Gamma=\Gamma_{0}H^{2}$. Upper Left Plot: This is the vector plot of the one dimensional system (32). Here, we adopt $\beta=3$ but for other values of $\beta>0$ similar type of plot can be found . Upper Right Plot: The phase plot of the system (34) when we have assumed $w=1.5$ and $\beta=2$. For other values of $w>0$ and $\beta>0$, we can also obtain similar phase space structure. Lower Left Plot: The phase space of the system (34) when the equation of state $w$ takes the value 00. Here we use $\beta=2$ but any positive value of $\beta$ gives same type of phase portrait. Lower Right Plot: The phase plot of the system (34) considering $w=-0.8$ and $\beta=3$. Also, we can get similar type of graphics for any positive value of $\beta$ and negative value of $w$ in the interval $(-1,0)$. Here the green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). In all three two-dimensional graphs, the black dotted curve separating the cyan and yellow regions corresponds to $q=-1$.
Figure 6: This figure also depicts the phase space of the two-fluid system driven by the matter creation rate $\Gamma=\Gamma_{0}H^{2}$ but here we consider the case when $w\leq-1$. Left Plot: The phase space of the system (34) when the equation of state $w$ takes the value $-1$. Here we use $\beta=3$ but any positive value of $\beta$ gives similar type of phase portrait. Right Plot: The phase plot of the system (34) when we assume $w=-1.5$ and $\beta=3$. For other values of $w<-1$ and $\beta>0$, we can also obtain similar graphics. Here the green region corresponds to the decelerating phase ($q>0$), the cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve (in the left graph this curve is not visible because of the presence of the critical line) separating the cyan and yellow regions corresponds to $q=-1$.
  1. 1.
@@
    When $w>0$: Considering the equation of state ($w$) of the second fluid as positive, one can get five isolated critical points, namely, $C_{0},~{}C_{1},~{}C_{2},~{}C_{3},~{}C_{4}$ which are highlighted in the Table 3. Now investigation of eigenvalues assure that $C_{1},~{}C_{3}$ are stable points and $C_{2},~{}C_{4}$ are saddle points. Again, eigenvectors corresponding to the eigenvalues $\beta,~{}0$ at the point $C_{0}$ lie on $x$ and $z$ axis respectively. Since, $x$ is increasing on $z$ axis and $z$ is increasing on $x$ axis, $C_{0}$ behaves like an unstable point. Here, $C_{1},~{}C_{4}$ belong to the accelerating phase and $C_{2},~{}C_{3}$ belong to the decelerating phase, while decelerating parameter is undefined at $C_{0}$. Now, the condition to be a viable trajectories are the ones that at $z=1/2$ the deceleration parameter should be negative. This implies $x\left(1-{\beta}\right)+(1+3w)(1-x)<0\Longrightarrow x>\frac{1+3w}{\beta+3w}.$ The upper right plot of the Fig. 5 depicts the complete evolution of the phase space orbits. Here, note that one can trace a set of orbits for which the universe leaves the past decelerating phase, enters in the present (or at finite time) accelerating epoch and ends again in a DM dominated decelerating phase or a DM dominated super accelerating phase.
  2. 2.
@@
    When $w=0$: If we assume the equation of state $w$ is zero, we obtain three isolated critical points $C_{0},~{}C_{1},~{}C_{4}$ which are mentioned earlier and one critical line $C_{7}(x_{c},1)$. Again, one can obtain $z^{\prime}>0$, $z^{\prime}<0$ and $x^{\prime}>0$ according to $z>\frac{\beta x}{3+\beta x}$, $z<\frac{\beta x}{3+\beta x}$ and on $z=0$ line, respectively if we choose $\beta$ as positive parameter. Therefore, the above arguments indicate that $C_{7}$ gives stable like nature, $C_{0}$ is unstable, $C_{1}$ is stable and $C_{4}$ is saddle point. Note that $C_{1},~{}C_{4}$ are accelerated points, but $C_{7}$ represents decelerating phase. The lower left plot of the Fig. 5 highlights the qualitative behavior of this model. It is also noted that the universe leaves its past decelerating phase and then it enters into the current (or at finite time) accelerating phase and finally it again enters into a decelerating phase admitting scaling solutions or a super accelerating phase dominated by DM.
  3. 3.
@@
    When $-\frac{1}{3}\leq w<0$: For these values of $w$, five isolated critical points ($C_{0},~{}C_{1},~{}C_{2},~{}C_{3},~{}C_{4}$) always belong to our domain $R_{3}$. From the sign of the eigenvalues, one can conclude that $C_{1},~{}C_{2}$ act like stable points and $C_{3},~{}C_{4}$ are saddle by nature. As before with similar reason, $C_{0}$ shows unstable character. Depending on the decelerating parameter, $q$, the critical points $C_{1},~{}C_{4}$ lie in the accelerating phase and $C_{2},~{}C_{3}$ also lie in the decelerating phase, while we can say nothing about the acceleration of the point $C_{0}$. The phase plot is similar to the lower right plot of the Fig. 5.
    Thus, the final state of the universe is second fluid dominated decelerating phase or DM dominated super accelerating epoch.
  4. 4.
@@
    When $-1<w<-\frac{1}{3}$: The qualitative behavior of the critical points given in the Table 3 are exactly similar to the above scenario where $w$ lies in $-\frac{1}{3}\leq w<0$, which are exhibited in the lower right plot of the Fig. 5. Note that, here the point $C_{2}$ represents acceleration. Here, the fate of the universe is completely DE dominated accelerating epoch or DM dominated super accelerating phase.
  5. 5.
@@
    When $w=-1$: In this case, the dynamical system (33) gives two isolated critical points $C_{1}$, $C_{3}$ which are described in Table 3, and two extra critical lines, namely, $C_{5}(0,z_{c})$ and $C_{6}\left(x_{c},\frac{\beta}{\beta+3}\right)$. There are no restriction on existence of the critical points $C_{1},~{}C_{3}$ and the critical line $C_{5}$ but the critical line $C_{6}$ belongs to our domain $R_{3}$ only when $\beta$ is taken to be positive. If we look into the evolution of $z$, one can easily find that $z^{\prime}$ is positive and negative when we choose $z>\frac{\beta}{\beta+3}$ and $z<\frac{\beta}{\beta+3}$, respectively. Again, $x$ is increasing below the line $z=\frac{\beta}{\beta+3}$ and it is decreasing above the line $z=\frac{\beta}{\beta+3}$. As a result, $C_{1}$ is stable, $C_{3}$ is always saddle, $C_{6}$ gives unstable behavior, $C_{5}$ is stable for $z_{c}>\frac{\beta}{\beta+3}$ otherwise this line is unstable. Clearly, the critical point $C_{1}$ and the critical line $C_{5},~{}C_{6}$ lead to the accelerated solution, while $C_{3}$ gives decelerated solution. The phase space structure are clearly visible in the left plot of Fig. 6. Therefore, $C_{5}$ with $z_{c}>\frac{\beta}{\beta+3}$ represents completely DE dominated accelerating solution.
  6. 6.
@@
    When $w<-1$: If we consider the second fluid as dark fluid whose equation of state lie in the phantom region, the critical points for the dynamical system (34) are $C_{0},~{}C_{1},~{}C_{2},~{}C_{3},~{}C_{4}$ which all are isolated. Inspection of the eigenvalues and the direction of the flow on the $x,~{}z$ axes, we can easily say that $C_{0},~{}C_{2},~{}C_{3}$ are saddle by behavior, $C_{4}$ is unstable and $C_{1}$ is the only globally stable point. Here, $C_{1},~{}C_{3},~{}C_{4}$ are matter dominated solutions and $C_{0},~{}C_{2}$ correspond to completely DE dominated solution. Again, $C_{1},~{}C_{2},~{}C_{4}$ are accelerated critical points and $C_{3}$ gives deceleration. The right plot of Fig. 6 shows the correct phase portrait. Although, we assume that the second fluid is phantom-like, however, we do not notice any DE dominated accelerating late time stable point, rather we find DM dominated super accelerating $(q<-1)$ late time stable point.

### v2

@@ III.3 Model: $\Gamma=\Gamma_{0}H^{2}$
In this case, we have constructed the model considering the particle creation rate $\Gamma=\Gamma_{0}H^{2}$ where $\Gamma_{0}$ is a constant having dimension of the inverse of the Hubble rate. We shall also discuss how $\Gamma$ influences the universe’s evolution.
@@
where $\beta=\Gamma_{0}H_{0}$ is a dimensionless and positive constant. To resolve the singularity at $z=0$ in the dynamical system (25), we redefine the time variable by setting $dN=zd\tau$. With this transformation, the reduced autonomous system, which remains topologically equivalent to system (25), takes the form:
$\frac{dx}{d\tau}=$ $x(1-x)\left(3wz+\beta(1-z)\right),$ (26a)
$\frac{dz}{d\tau}=$ $\frac{3}{2}z(1-z)\left[\left(1+w(1-x)\right)z-\frac{\beta x(1-z)}{3}\right].$ (26b)
Now, for the dynamical system (26), it follows that x=0,x=1,z=0,z=1x=0,~x=1,~z=0,~z=1 are invariant manifolds, which leads to the fact that the domain $\bf R$} is positively invariant. Being $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\beta x(1-z)}{z}\right]$ in this case, the requirement for realizing an accelerating phase at the present moment (i.e. $z=1/2$), the physical trajectories should satisfy the relation $(1-\beta)x+(1+3w)(1-x)<0$. In the following we describe the qualitative features of the above dynamical system for different values of $w$.
Critical point Existence Eigenvalue Stability $\mathbf{\Omega_{f}}$ $\mathbf{\Omega_{\rm dm}^{\rm eff}}$ $\mathbf{q}$ Acceleration
$C_{0}(0,0)$ Always $(\beta,0)$ Unstable if w>1w>-1; 1 0 Undefined Undetermined
Non-hyperbolic Saddle if w<1w<-1
$C_{1}(1,0)$ Always $\left(-\beta,-\frac{\beta}{2}\right)$ Stable 0 1 -$\infty$ Yes
$C_{2}(0,1)$ Always $\left(3w,-\frac{3}{2}(1+w)\right)$ Stable if 1<w<0-1<w<0; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if w<1w<-1 and if $w>0$
@@
$C_{4}\left(1,\frac{\beta}{\beta+3}\right)$ $\beta>0$ $\left(-\frac{3\beta(1+w)}{3+\beta},\frac{3\beta}{2(3+\beta)}\right)$ Saddle if w>1w>-1 0 1 1-1 Yes
Unstable if w<1w<-1
$C_{5}\left(0,z_{c}\right)$ w=1w=-1 $\left(\beta-(\beta+3)z_{c},0\right)$ Stable if $z_{c}>\frac{\beta}{\beta+3}$ 11 0 1-1 Yes
@@
$C_{6}\left(x_{c},\frac{\beta}{\beta+3}\right)$ w=1w=-1 $\left(0,\frac{3\beta x_{c}}{2(3+\beta)}\right)$ Unstable $1-x_{c}$ $x_{c}$ 1-1 Yes
@@
Table 3: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (26) with the matter creation rate $\Gamma=\Gamma_{0}H^{2}$.
@@
Figure 7: Description of the phase space for the matter creation rate $\Gamma=\Gamma_{0}H^{2}$. Upper Left Plot: The phase plot of the system (26) when we have assumed $w=0.1$ and $\beta=2$. For other values of $w>0$ and $\beta>0$, we can also obtain similar phase space structure. Upper Middle Plot: The phase space of the system (26) when the EoS $w$ takes the value 0. Here we use $\beta=2$ but any positive value of $\beta$ gives same type of phase portrait. Upper Right Plot: The phase plot of the system (26) considering w=0.8w=-0.8 and $\beta=3$. Also, we can get similar type of graphics for any positive value of $\beta$ and negative value of $w$ in the interval (1,0)(-1,0). Lower Left Plot: The phase space of the system (26) when the EoS $w$ takes the value 1-1. Here we use $\beta=3$ but any positive value of $\beta$ gives similar type of phase portrait. Lower Right Plot: The phase plot of the system (26) when we assume w=1.5w=-1.5 and $\beta=3$. For other values of w<1w<-1 and $\beta>0$, we can also obtain similar graphics. Here the green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with 1<q<0-1<q<0 and the yellow region corresponds to the super accelerating phase (i.e. q<1q<-1). In all five graphs, the black dotted curve (in the lower left graph this curve is not visible because of the presence of the critical line) separating the cyan and yellow regions corresponds to q=1q=-1.
Figure 8: We have presented the evolution of key cosmological parameters namely, the DM density parameter $\Omega_{\rm dm}^{\rm eff}$, the second fluid density parameter $\Omega_{f}$, the decelerating parameter $q$, and the effective equation of state of DM $w_{\rm dm}^{\rm eff}$ for the dynamical system (26) with the matter creation rate $\Gamma=\Gamma_{0}H^{2}$. Upper Left Plot: This plot is appeared for the model parameter $w=0.1$ and $\beta=1$. Numerical simulation shows that the final fate of the universe is attracted by DM dominated decelerated phase. Upper Right Plot: This plot is also appeared for $w=0.1$, $\beta=1$, but the trajectories finish in a DM dominated phantom phase. Lower Left Plot: For this plot, we have taken the model parameters values $w=0$ and $\beta=1$. Numerical simulation depicts that the ultimate fate of the universe is attracted in a decelerated phase. Lower Right Plot: For this plot, we have also taken the model parameters values $w=0$, $\beta=1$, but the trajectories finish in a DM dominated phantom phase. In these scenarios, no decelerating DM-dominated saddle point is found (see upper left and middle plots of Fig. 7), and hence the past DM-dominated phase can not be realized. Again, each orbit originates from the critical point $C_{0}$, where $w_{\rm dm}^{\rm eff}$ diverges to -$\infty$. Consequently, $w_{\rm dm}^{\rm eff}$ diverges during the early phase.
[…]
@@
When $w>0$, one can get five isolated critical points, namely, $C_{0},~C_{1},~C_{2},~C_{3},~C_{4}$ which are highlighted in the Table 3. Now investigation of eigenvalues assure that $C_{1},~C_{3}$ are stable points and $C_{2},~C_{4}$ are saddle points. Again, eigenvectors corresponding to the eigenvalues $\beta,~0$ at the point $C_{0}$ lie on $x$ and $z$ axis respectively. Since, $x$ is increasing on $z$ axis and $z$ is increasing on $x$ axis, $C_{0}$ behaves like an unstable point. Here, $C_{1},~C_{4}$ belong to the accelerating phase and $C_{2},~C_{3}$ belong to the decelerating phase, while decelerating parameter is undefined at $C_{0}$. Now, the condition to be a viable trajectories are the ones that at $z=1/2$ the deceleration parameter should be negative. This implies $x\left(1-{\beta}\right)+(1+3w)(1-x)<0\Longrightarrow x>\frac{1+3w}{\beta+3w}.$ The upper left plot of the Fig. 7 depicts the complete evolution of the phase space orbits. Here, note that one can trace a set of orbits for which the universe leaves the past decelerating phase, enters in the present (or at finite time) accelerating epoch and ends again in a DM dominated decelerating phase or a DM dominated super accelerating phase, which we have demonstrated in the upper left and right plots in Fig. 8. As seen in this plot, there is no past DM dominated epoch, and the evolution of $w_{\rm dm}^{\rm eff}$ which diverges in the early phase.
@@
When $w=0$, we obtain three isolated critical points $C_{0},~C_{1},~C_{4}$ which are mentioned earlier and one critical line $C_{7}(x_{c},1)$. Again, one can obtain $z^{\prime}>0$, $z^{\prime}<0$ and $x^{\prime}>0$ according to $z>\frac{\beta x}{3+\beta x}$, $z<\frac{\beta x}{3+\beta x}$ and on $z=0$ line, respectively if we choose $\beta$ as positive parameter. Therefore, the above arguments indicate that $C_{7}$ gives stable like nature, $C_{0}$ is unstable, $C_{1}$ is stable and $C_{4}$ is saddle point. Note that $C_{1},~C_{4}$ are accelerated points, but $C_{7}$ represents decelerating phase. The upper middle plot of the Fig. 7 highlights the qualitative behavior of this model. It is also noted that the universe leaves its past decelerating phase and then it enters into the current (or at finite time) accelerating phase and finally it again enters into a decelerating phase admitting scaling solutions or a super accelerating phase dominated by DM. This scenario is clearly noticed in the lower left and right plots in the Fig. 8. It is evident from this plot that no past DM dominated era occurs, and $w_{\rm dm}^{\rm eff}$ becomes divergent in the early phase.
@@
When $\frac{1}{3}\leq w<0$, five isolated critical points ($C_{0},~C_{1},~C_{2},~C_{3},~C_{4}$) always belong to our domain $R$. From the sign of the eigenvalues, one can conclude that $C_{1},~C_{2}$ act like stable points and $C_{3},~C_{4}$ are saddle by nature. As before with similar reason, $C_{0}$ shows unstable character. Depending on the decelerating parameter, $q$, the critical points $C_{1},~C_{4}$ lie in the accelerating phase and $C_{2},~C_{3}$ also lie in the decelerating phase, while we can say nothing about the acceleration of the point $C_{0}$. The phase plot is similar to the upper right plot of the Fig. 7. Thus, the final state of the universe is second fluid dominated decelerating phase or DM dominated super accelerating epoch.
@@
When $1<w<-\frac{1}{3}$, the qualitative behavior of the critical points given in the Table 3 are exactly similar to the above scenario where $w$ lies in $\frac{1}{3}\leq w<0$, which are exhibited in the upper right plot of the Fig. 7. Note that, here the point $C_{2}$ represents acceleration. Here, the fate of the universe is completely DE dominated accelerating epoch or DM dominated super accelerating phase.
@@
When w=1w=-1, the dynamical system (25) gives two isolated critical points $C_{1}$, $C_{3}$ which are described in Table 3, and two extra critical lines, namely, $C_{5}(0,z_{c})$ and $C_{6}\left(x_{c},\frac{\beta}{\beta+3}\right)$. There are no restriction on existence of the critical points $C_{1},~C_{3}$ and the critical line $C_{5}$ but the critical line $C_{6}$ belongs to our domain $R$ only when $\beta$ is taken to be positive. If we look into the evolution of $z$, one can easily find that $z^{\prime}$ is positive and negative when we choose $z>\frac{\beta}{\beta+3}$ and $z<\frac{\beta}{\beta+3}$, respectively. Again, $x$ is increasing below the line $z=\frac{\beta}{\beta+3}$ and it is decreasing above the line $z=\frac{\beta}{\beta+3}$. As a result, $C_{1}$ is stable, $C_{3}$ is always saddle, $C_{6}$ gives unstable behavior, $C_{5}$ is stable for $z_{c}>\frac{\beta}{\beta+3}$ otherwise this line is unstable. Clearly, the critical point $C_{1}$ and the critical line $C_{5},~C_{6}$ lead to the accelerated solution, while $C_{3}$ gives decelerated solution. The phase space structure are clearly visible in the lower left plot of Fig. 7. Therefore, $C_{5}$ with $z_{c}>\frac{\beta}{\beta+3}$ represents completely DE dominated accelerating solution.
@@
When w<1w<-1, the critical points for the dynamical system (26) are $C_{0},~C_{1},~C_{2},~C_{3},~C_{4}$ which all are isolated. Inspection of the eigenvalues and the direction of the flow on the x,zx,~z axes, we can easily say that $C_{0},~C_{2},~C_{3}$ are saddle by behavior, $C_{4}$ is unstable and $C_{1}$ is the only globally stable point. Here, $C_{1},~C_{3},~C_{4}$ are matter dominated solutions and $C_{0},~C_{2}$ correspond to completely DE dominated solution. Again, $C_{1},~C_{2},~C_{4}$ are accelerated critical points and $C_{3}$ gives deceleration. The lower right plot of Fig. 7 shows the correct phase portrait. Although, we assume that the second fluid is phantom-like, however, we do not notice any DE dominated accelerating late time stable point, rather we find DM dominated super accelerating (q<1)(q<-1) late time stable point.
In Fig. 9, the normalized Hubble parameter $H/H_{0}$ is plotted for the matter creation model ($\Gamma=\Gamma_{0}H^{2}$) alongside $\Lambda{\rm CDM}$ (upper panel). The fractional deviation $\Delta H/H$ is shown below. We observe that the two models agree at low redshift, but at higher-redshift, deviations reveal the influence of particle creation.